OpenStudy (anonymous):

Can anyone explain how you can determine how much power a circuit will generate if you know the voltage and resistance but not the current?

OpenStudy (jamesj):

The power is \[ P = VI \] Now because \( V = IR \), that means we can also write power as \[ P = I^2R \] or \[ P = V^2/R \]